OpenStudy (anonymous):

Marcus wants to purchase a home in six years. He will contribute $2500 each year to a savings account with 2.64% interest, compounded semiannually. What is the future value of this investment, when Maurice needs to make a down payment?

OpenStudy (anonymous):

@Lena772

OpenStudy (lena772):

I don't know, sorry.

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

do you have a financial calculator? if so just plug in N = 12 I = 2.64/2 PV = 0 PMT = 2500 compute FV

OpenStudy (dumbcow):

oh wait sorry that is not right first find effective annual growth rate \[(1 + \frac{.0264}{2})^2 - 1\] set this as I N = 6 PV =0 PMT = 2500

OpenStudy (anonymous):

whats next

OpenStudy (dumbcow):

thats it, it will compute the future value do you have a financial calculator?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Katherine invested $1225 in a savings account which earns 3.25% interest compounded semiannually. What will the account be worth in 7 years?

OpenStudy (dumbcow):

haha what , that is same question as your last post i gave the formula for compound interest to solve this problem

OpenStudy (dumbcow):

so you have to do it by hand....hmm im guessing your class gave you a list of formulas to use then

OpenStudy (dumbcow):

http://en.wikipedia.org/wiki/Future_value#Compound_interest \[FV = 2500* \frac{(1+r)^n - 1}{r}\] n = 6 r = .02657424